OpenStudy (anonymous):

how do i find the minimum and maximum possible areas of a rectangle measuring 9m by 5m

OpenStudy (anonymous):

can someone please help me with a few math questions, thanks

OpenStudy (anonymous):

Francesca reads 57 pages every 2 h. How many pages can she read in 5 h?

OpenStudy (gorv):

welcome to OS both of u :)

OpenStudy (gorv):

@lyd wecan not find min or max area as we can not change the length of sides of rectangle

OpenStudy (gorv):

its area=9*5=45 fix it is

OpenStudy (gorv):

@harris6607 dont spoil others Q buddy ..ask your Q seprately ...that will good bcoz some one will help u for sure

OpenStudy (gorv):

ok in 2 hr she read = 57 page in 1 hr she read=57/2 in 5 hr she will read = (57/2)*5